Nitrogen (N), Phosphorus (P), and Potassium (K) are the main nutrients in plant fertilizers. According to an industry convention, the numbers on the label refer to the mass percents of N, P2O5, and K2O, in that order. Calculate the N:P:K ratio of a 30:10:10 fertilizer in terms of moles of each element, and express it as X:y:1.0. (Assume 100 g)
1
Expert's answer
2016-09-20T05:23:19-0400
Answer:
Molar mass of N is 14.0067 g/mol.
Molar mass of P2O5 is 30.973762*205 = 6349.62121 g/mol.
Molar mass of K2O is 39.0983*20 = 781.966 g/mol/

Ratio of masses is 30:10:10
Therefore for each 50 g of fertilizer:
For N: 30 g / 14.0067 g/mol = 30/14.0067 moles
For P2O5: 10 g / 6349.62121 g/mol = 10/6349.62121 moles
For K2O: 10 g / 781.966 g/mol = 10/781.966 moles

Ratio of elements in terms of moles:
30/14.0067 : 10/6349.62121 : 10/781.966
30/14.0067*781.966/10 : 10/6349.62121*781.966/10 : 10/781.966*781.966/10
167.484 : 0.123 : 1.0